OpenStudy (anonymous):

When f(x) is divided by h(x)=x+2, q(x)=2(x^2)-x+2 and r(x)=-2. Find f(x).

OpenStudy (anonymous):

2x^2+2=q(x)+x

OpenStudy (anonymous):

f(x)=h(x)*q(x)+r(x)

OpenStudy (anonymous):

could you explain how you got that?

OpenStudy (anonymous):

Actually i think q(x) stands fr quotient and r(x) for remainder.

OpenStudy (anonymous):

oh! ok, it makes more sense now.